# PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling

### v1

@@ 1 Introduction
Open-ended Generation ? $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$ $\times$ $\times$ $\times$
Conditional Generation ? $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Table 1: Categorization of NLP tasks based on constraint aspects. $\checkmark$ indicates full support, $\times$ indicates no support, and $\triangle$ indicates partial or limited support.
As diffusion models significantly enhance the quality and diversity of generated outputs in continuous domains such as images and audio (Song et al., 2021b), recent research has increasingly applied diffusion models to NLP  (Li et al., 2022; Gong et al., 2023a; Yuan et al., 2023; He et al., 2023).
To integrate the diffusion process with traditional language models, research has predominantly followed two approaches: Continuous Diffusion Language Model (CDLM) (Yuan et al., 2023; Lovelace et al., 2023; Chen et al., 2023), which embed text in a continuous latent space, and Discrete Diffusion Language Model (DDLM) (He et al., 2023; Lou et al., 2024; Zhou et al., 2024; Shi et al., 2024; Sahoo et al., 2024; Zheng et al., 2024), which operate directly on the vocab space. Recent DDLMs adeptly reflect the highly structured nature of language; this property enables sequence control in natural language, generally resulting in high performance in NLP fields such as unconditional generation and open-ended generation  (Lou et al., 2024). However, for DDLMs to be widely adopted across various NLP domains, particularly in tasks such as close-ended conditional generation — hereafter referred to as dataset-guided generation — existing DDLMs still exhibit shortcomings in their performance. Our experiments suggest that models such as SEDD falls short in such tasks, potentially limiting their practical applications.
@@
Better performance can be achieved by integrating Pretrained Language Models (PLMs) through the use of an encoder-based PLM as the initial setup for embedding functions, leveraging its capability to capture overall semantics (He et al., 2023). However, integrating PLMs into diffusion models is non-trivial as PLMs typically predict masked elements in a single step, whereas diffusion models require step-wise denoising based on the overall semantics of each timestep sequence, and such gap yields limited results. Therefore, we need to consider such inconsistency to effectively adopt PLMs into DDLMs by a new methodology.
@@
In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates Mask Language Model (MLM)-based PLMs with DDLMs for dataset-guided generation. To address the gap between the step-wise nature of diffusion models and the one-step prediction strategy of PLMs, we begin by considering MLM-based PLMs as the denoising function at each step of the diffusion process. Since identifying where to apply denoising during the diffusion process is challenging for PLMs, we incorporate an entropy tracking module to support their operation. For constructing entropy tracking module, we interpret each denoising step as a constrained Markov Random Field (cMRF), which enables us to adopt adaptive Gibbs sampling, which then facilitates leveraging entropy information from conditional sequences. To train DDLMs to effectively exploit an entropy tracking module, we propose a forward process based on an entropy-based noising scheduling during training phase. Specifically, our entropy-based noising scheduling noises lower entropy tokens first, thereby learning to progressively generate sequences guided by the entropy from entropy tracking module, which enhances the quality of the generated text.
@@
Experimental results demonstrate that Diffusion-EAGS achieves outstanding performance compared to existing autoregressive model (GPT-2), CDLMs, and DDLMs across various dataset-guided generation tasks. We demonstrate that both PLMs and entropy-based Gibbs sampling contribute to higher performance in our ablation study. Furthermore, our model exhibits higher diversity in certain tasks compared to LLMs and can facilitate token-level generation, indicating the potential applicability of our model across a wide range of dataset-guided generation tasks. In addition, we validate that our model can adapt to both bilingual and low-resource settings, which are frequently encountered in practical applications.
@@ 2 Task Setting
2.1 Dataset-Guided Generation
In sequence-to-sequence (seq2seq) conditional generation tasks, the level of constraint imposed by the dataset plays a critical role in shaping the generation process. As shown in Table 1, constraints can be categorized into three levels: the provision of context alone, the provision of specific content that should be covered in the target sequence, and the provision of content format. Each level represents an increasing degree of constraint in the task requirements. The goal is to develop a model capable of generalizing across a wide range of downstream tasks, rather than being specialized for a specific task, while adhering to the constraints of the given dataset—a task we refer to as Dataset-Guided Generation.
@@ 2.2 DDLM with BERT
Diffusion Models The process of diffusion models consists of two main stages: the forward process, where noise is added to the input over multiple steps, and the reverse process, where noise is removed to restore the information. There are two types: Continuous and Discrete diffusion models. Discrete diffusion models operate the diffusion process across categorical random variables Hoogeboom et al. (2021). As discrete diffusion models show promising results in recent studies Venkatraman et al. (2024); Sahoo et al. (2024); Lou et al. (2024), we set up our language models grounded on discrete diffusion language models.
BERT as constrained Markov Random Field (cMRF) As the pretraining paradigm has significantly contributed to the success of language models, the integration of PLMs is essential in the development of language models. Nevertheless, adopting PLMs to DDLMs is a complex task, as DDLMs require a step-wise denoising strategy, while MLM-based PLMs are trained to predict masked elements in a single step. To effectively integrate PLMs into DDLMs, it is necessary to resolve such a discrepancy. Therefore, we begin by interpreting the intrinsic characteristics of MLMs as Markov Random Field (MRF), as demonstrated in Wang and Cho (2019). Details are in Appendix A.
Figure 1: Potential Function Difference of MLM based pre-trained BERT generation and RocStories dataset-guided BERT generation
However, dataset guided generation occupy different fields in terms of randomness, compared to open-ended generation. In Figure 1, we can observe an increase of potentials for same sequence X("The quick brown fox jumps over the lazy dog.") on RocStories. That is, it is necessary to investigate whether the dataset-constrained space is still subject to MRF.
Given a sequence of fully connected random variables $X=(x_{1},x_{2},\ldots,x_{L})$ with a set of observed variables $Y=(y_{1},y_{2},\ldots,y_{K})$, $f_{\theta}(X\backslash\{x_{l}\};Y)$ can be represented by logits from MLM given the sequence $X\backslash\{x_{l}\}$ and the observed variables $Y$. Then, similarly to Equation 2, log-potential can be calculated as:
$\log\phi_{l}(X;Y)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y),&\text{if }\text{[M]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$
By this, we can define the potential function for the clique as:
$\phi(X;Y)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X;Y)\right)$
@@ 3 Methodology
Figure 2: Overview of Diffusion-EAGS. During forward steps, We schedule the noise based on ENS using PLM. In generation steps, we denoise the sequences based on EAGS.
We propose Diffusion-EAGS that integrates PLM with DDLMs in dataset-guided generation, through Entropy-based Adaptive Gibbs Sampling (EAGS) and Entropy-based Noise Scheduling (ENS). The overview of our process is in Figure 2.
@@ 3.1 Generation Process: Entropy-based Adaptive Gibbs Sampling (EAGS)
Building on the premise in Section 2.2, we propose EAGS to leverage PLM as the function of denoising process in diffusion models. As MLM-based PLMs can be interpreted as cMRF, we can apply Gibbs sampling in each denoising step in DDLM. By Gibbs sampling, we can sample each token $x_{l}$ from its constrained distribution $p(x_{l}|X\backslash\{x_{l}\};Y)$. Rather than performing Gibbs sampling, we employ EAGS, as the entropy of each variable indicates the amount of information at certain position in the sequence. By generating tokens in the order of decreasing entropy, we prioritize the generation of the least informative parts of the sequence. Such prioritized generation makes a more structured sequences, leveraging the already established syntactic context.
According to Equation 1, we can compute the conditional probability for each masked position. The entropy $H$ for a given token $x_{l}$ is derived as:
$\begin{split}&H(x_{l}|Y,f)=\\ &-\sum_{m=1}^{M}p(x_{l}=v_{m}|x_{l},Y,f)\log p(x_{l}=v_{m}|x_{l},Y,f)\end{split}$
@@
    Entropy Calculation: Compute the entropy $H(x_{l}\mid Y,f)$ for each variable $x_{l}$.
@@
    Sampling: Sample $x_{l^{*}}$ from its conditional distribution $p(x_{l^{*}}\mid Y,f^{*})$.
@@ 3.2 Forward Process: Entropy-based Noise Scheduling (ENS)
To enhance the facilitation of Adaptive Gibbs Sampling in the generation process, it is essential to describe a similar process in the training phase. Therefore, we schedule the forward process of diffusion training based on the entropy $H(x_{l})$ of positions $x_{l}$ with the input sequence $[Y|X]$. Specifically, the position-wise entropy is calculated by the PLM that shares a similar language base with diffusion-trained model. Assuming the diffusion process progresses over $T$ steps, at each step $t$, we mask the $L/T$ number of positions with the lowest entropy from the set $\{x_{1},\ldots,x_{L}\}$. The masking process at step $t$ in position $l$ is described by the denoising matrix $Q_{tl}$.
@@
This entropy-based noise scheduling approach ensures that the forward process in diffusion training closely mirrors the generation process, thereby enhancing the effectiveness of Adaptive Gibbs Sampling in language generation.
3.3 Training & Inference
Training Process
@@ 3.3 Training & Inference
for $BatchStep=0..N$  do
       $x\sim D$
       $\text{t}\sim\text{Randint}(0,T)$
       $f\leftarrow PLM(x|Y)$
       $\mathcal{H}\leftarrow H(x|Y,f)$
       $x^{t}\leftarrow\text{Forward}(x_{0},\mathcal{H},t)$
       $x^{t+1}\leftarrow\text{Forward}(x_{0},\mathcal{H},t+1)$
       $L_{s}=-\sum_{i}(q(x^{t}_{i})\mid x_{t+1})\log(p_{\theta}(x^{t}_{i})\mid x_{t+1})$
@@
       if  $t=T$  then // Initialize $x^{T}$
             $x^{T}\leftarrow G_{M}(L,Y)$
       else  // Denoise step
             $f\leftarrow M(x^{t},Y)$
[…]
             $x^{t-1}=\text{Sampling}(x^{t},l^{*},Y)$
@@
Algorithm 1 Training & Generation
Our whole process is in Algorithm 1. Distinct from the prevailing methodologies in diffusion models as described by Ho et al. (2020a) and Austin et al. (2023), we do not employ the PLM parameterization strategy $\widetilde{p}_{\theta}(\widetilde{z}_{0}|z_{t},t)$, which preserves the original semantic embedding spaces during the training phase. We empirically find that PLM parameterization restricts the diversity of generated responses. Instead, to ensure the completeness of sentences, we implement Minimum Bayes Risk (MBR) decoding in the final generation step. We follow the traditional diffusion loss in Equation (Ho et al., 2020b) with changing MSE to Cross Entrpy Loss.
@@ 4.1 Tasks and Datasets
Question Generation We use the Quasar-T dataset (Dhingra et al., 2017) to generate questions from passages and answers. Paraphrase The QQP dataset (Wang et al., 2017) assesses paraphrase detection by treating one question as a paraphrase of another. Social Datasets The Paradetox dataset (Logacheva et al., 2022) focuses on removing profanities. The Deontology from ETHICS dataset (Hendrycks et al., 2023) evaluates ethical judgment based on scenarios. Open-ended Generation The RocStories (Mostafazadeh et al., 2016) dataset is used to generate coherent story continuations from a given context. Detailed explanations are available in Appendix B and Dataset statistics are shown in Table 8.
@@ 4.2 Baselines
Since our objective is centered on generation tasks, we compare Diffusion-EAGS with three categories of baselines with the model similar to the size of RoBERTa-base: AutoRegressive (AR) model, CDLM, and DDLM. For AR model (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019), which is renowned for its proficiency in generation tasks. For CDLMs, our baseline includes DiffuSeq (Gong et al., 2023a) and DINOISER (Ye et al., 2024). DiffuSeq is a diffusion model specifically designed for sequence-to-sequence applications, whereas DINOISER adaptively determines the range of sampled noise scales during training. For DDLMs, we utilize DiffusionBERT (He et al., 2022) , a state-of-the-art model in the research series of DiffusionLM (Li et al., 2022), and SEDD (Lou et al., 2024), a state-of-the-art model of open-ended generation in diffusion language generation domain. For SEDD, we download the pretrained version and fine-tune on it. Experimental results on LLMs can be found in Appendix C.

### v2

@@ 1 Introduction
Open-ended Generation $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$ $\times$ $\times$ $\times$
Conditional Generation $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Table 1: Each dataset has a different level of conditional contraints even if they are conditional generation tasks. $\checkmark$ indicates full support, $\times$ indicates no support, and $\triangle$ indicates partial or limited support.
As diffusion models significantly enhance the quality and diversity of generated outputs in continuous domains such as images and audio (Song et al., 2021b), recent research has increasingly applied diffusion models to NLP, and demonstrate their promising performance over autoregressive models (Li et al., 2022; Gong et al., 2023a; He et al., 2023; Yuan et al., 2023; Lovelace et al., 2023; Chen et al., 2023; He et al., 2023; Lou et al., 2024; Zhou et al., 2024; Shi et al., 2024; Sahoo et al., 2024; Zheng et al., 2024; Nie et al., 2024). Moreover, these models have been shown to be scalable, and are able to address the challenge of generalizing bidirectional relationships while maintaining sensitivity to the temporal alignment between training and test data, a conventional architectural limitation often observed in autoregressive models (Nie et al., 2024).
Although recent Discrete Diffusion Language Models (DDLMs) have demonstrated high performance in NLP tasks such as unconditional and open-ended generation compared to Continuous Diffusion Language Models (CDLMs) (Lou et al., 2024), existing DDLMs continue to face limitations in conditional generation, thereby restricting their broader applicability across diverse NLP domains. Our experiments suggest that models such as SEDD falls short in such tasks.
@@
To enhance performance of DDLMs, PLMs can be integrated into diffusion models. This approach leverages the capability of PLMs to elicit improved generalization capabilities in unseen tasks while simultaneously exploiting the controllability and diversity strengths inherent to diffusion models. However, integrating PLMs into diffusion models is non-trivial as PLMs typically predict masked elements in a single step, whereas diffusion models require step-wise denoising based on the overall semantics of each timestep sequence, and such gap yields limited results. Therefore, we need to consider such inconsistencies to effectively adopt PLMs into DDLMs by a new methodology.
@@
In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates Mask Language Model (MLM)-based PLMs with DDLMs for conditional generation. To address the gap between the step-wise nature of diffusion models and the one-step prediction strategy of PLMs, we begin by adopting PLMs as the denoising function at each step of the diffusion process. Since identifying where to apply denoising during the diffusion process is challenging for PLMs in our experiments, we incorporate an entropy tracking module to support their operation based on entropy information. To train DDLMs to effectively exploit an entropy tracking module, we propose an entropy-based noising scheduling during training phase in a forward process. Specifically, our entropy-based noising scheduling noises lower entropy tokens first, thereby learning to progressively generate sequences guided by the entropy from entropy tracking module.
@@
Experimental results demonstrate that Diffusion-EAGS achieves outstanding performance compared to baseline models across various conditional generation tasks. Furthermore, our model exhibits higher diversity in certain tasks compared to LLMs and can facilitate token-level generation, and validate that our model can also adapt to both bilingual and low-resource settings, which are frequently encountered in practical applications, indicating the potential applicability of our model across a wide range of conditional generation tasks.
@@ 2 Task Setting
2.1 Fine-Grained Conditional Generation
In conditional generation tasks, the level of conditional constraint imposed by the dataset plays a critical role in shaping the generation process. As shown in Table 1, conditional constraints are diverse across datasets. In our task, we categorize these constraints into three levels: (1) the provision of context alone, requiring the continuity of the prefix; (2) the provision of specific content to be included in the target sequence, necessitating the inclusion of certain keywords; and (3) the provision of semantic content formatting, such as transforming toxic sentences into safer alternatives or converting text from the source language to a target language. In our study, we aim to develop a universal diffusion framework capable of being applied across a wide range of conditional generation tasks.
@@ 2.2 DDLM with BERT
BERT as Markov Random Field (MRF) As the pretraining paradigm has significantly contributed to the success of language models, the integration of PLMs is essential in the development of language models. Nevertheless, adopting PLMs to DDLMs is a complex task, as DDLMs require a step-wise denoising strategy, while PLMs are trained to predict masked elements in a single step. To effectively integrate PLMs into DDLMs, it is necessary to resolve such a discrepancy. Therefore, we begin by interpreting the intrinsic characteristics of MLMs as Markov Random Field (MRF), as demonstrated in Wang and Cho (2019). Details are in Appendix B.
constrained MRF (cMRF) Even though open-ended MLMs are MRFs, we observe a significant increase in log-potential values for sequences when guided by the RocStories conditions, as shown in Figure 1. Additional experiments supporting this observation are detailed in Appendix A.
Figure 1: When a condition is provided, the distribution of potential values for the samples is shifted on a logarithmic scale.
This suggests that conditional generation models differ from open-ended models in randomness, making it crucial to investigate whether conditionally constrained spaces adhere to MRF properties. Through the derivation process outlined in Appendix B, we demonstrate that these spaces retain the characteristics of an MRF.
Given a sequence of fully connected random variables $X=(x_{1},x_{2},\ldots,x_{L})$ with a set of observed variables $Y=(y_{1},y_{2},\ldots,y_{K})$, $f_{\theta}(X\backslash\{x_{l}\};Y)$ can be represented by logits from MLM given the sequence $X\backslash\{x_{l}\}$ and the observed variables $Y$. Then, log-potential can be calculated as:
$\log\phi_{l}(X;Y)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y),&\text{if }\text{[MASK]}\notin X_{l-1:l+1},\\ \\ 0,&\text{otherwise}\end{cases}$
where $1h(x_{l})$ is a one-hot vector with $x_{l}$ set to 1. By this, we can define the potential function for the clique as:
$\phi(X;Y)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X;Y)\right)$ (1)
@@ 3 Methodology
Figure 2: Given a conditional constraint $Y$ and a generation target $X$, the training process utilizes Entropy-based Sampling (ENS) to determine which positions in the masked sequence should be denoised when transitioning from one timestep to the next. The tokens at the corresponding masked positions ([M] denotes masked tokens) are then generated using $f_{\theta}$. The loss is subsequently computed, and the parameters are updated accordingly. During the inference process, the sequence $X$, initially composed of mask tokens, is iteratively refined through the diffusion process to generate $X$ that satisfies the given condition $Y$.
In this section, we propose Diffusion-EAGS that integrates PLM into DDLMs through Entropy-based Adaptive Gibbs Sampling (EAGS) and Entropy-based Noise Scheduling (ENS). The overview of our process is in Figure 2.
@@ 3.1 Generation Process: Entropy-based Adaptive Gibbs Sampling (EAGS)
As discussed in Section 2.2, MLM-based PLMs can be interpreted as cMRFs, allowing Gibbs sampling to be applied at each denoising step in DDLMs by leveraging the properties of MRFs. As PLMs often struggle to select the next denoising targets, it is important to highlight that the primary purpose of adopting Gibbs sampling is not only to bridge the gap between PLMs and diffusion models but also to assist PLMs in identifying which absorbing tokens should be denoised. Therefore, we propose Entropy-Adaptive Gibbs Sampling (EAGS) to address the challenge PLMs face in identifying the appropriate denoising targets. In EAGS, the entropy of each variable quantifies the amount of uncertainty at a specific position within the sequence. This enables sampling of each token $x_{l}$ from its constrained distribution $p(x_{l}|X\backslash\{x_{l}\};Y)$ in the descending order of entropy. By prioritizing the generation of the least informative parts of the sequence, EAGS facilitates the creation of more structured sequences by leveraging the syntactic context that has already been established.
Specifically, according to Equation (2), the probability $p^{t}(x_{l}=v_{m}|Y,f_{\theta})$ of the token $v_{m}$ given the context $Y$ and $f_{\theta}$ at step $t$ can be calculated as:
$p^{t}(x_{l}=v_{m}|Y,f_{\theta})=\frac{\exp(f_{\theta}(v_{m}|X^{t}))}{\sum_{m^{\prime}=1}^{M}\exp(f_{\theta}(v_{m^{\prime}}|X^{t}))}$
, where $f(v_{m}|X^{t})$ is the logit corresponding to the token $v_{m}$ for the position $x_{l}$ at step $t$. That is, we can compute the conditional probability for each masked position. With the conditional probability, the entropy $H$ for a variable $x_{l}$, logits function $f_{\theta}$, and the dataset context $Y$ is derived as:
$\begin{split}&H(x_{l}|Y,f_{\theta})=\\ &-\sum_{m=1}^{M}p(x_{l}=v_{m}|x_{l},Y,f_{\theta})\log p(x_{l}=v_{m}|x_{l},Y,f_{\theta})\end{split}$
@@
    Entropy Calculation: Compute the entropy $H(x_{l}\mid Y,f_{\theta})$ for each variable $x_{l}$.
@@
    Sampling: Sample $x_{l^{*}}$ from its conditional distribution $p(x_{l^{*}}\mid Y,f_{\theta}^{*})$.
@@ 3.2 Forward Process: Entropy-based Noise Scheduling (ENS)
To enhance the facilitation of Adaptive Gibbs Sampling in the generation process, it is essential to simulate a similar process in the training phase. Therefore, we schedule the forward process of diffusion training based on the entropy $H(x_{l})$ of position $x_{l}$ with the input sequence $[Y|X]$. Specifically, the position-wise entropy is calculated by the PLM that shares a similar language base with diffusion-trained model. Assuming the diffusion process progresses over $T$ steps, at each step $t$, we mask the $L/T$ number of positions with the lowest entropy from the set $\{x_{1},\ldots,x_{L}\}$. The masking process at step $t$ in position $l$ is described by the denoising matrix $Q_{tl}$.
@@
This entropy-based noise scheduling approach ensures that the forward process in diffusion training closely mirrors the generation process, thereby enhancing the effectiveness of EAGS in language generation.
3.3 Training and Inference
ENS Process:
@@ 3.3 Training and Inference
[…]
       $x\sim D$ // Sample data from $D$
       $t\sim\text{Randint}(0,T)$ // Sample random timestep
       $f\leftarrow\text{PLM}(x\mid Y)$ // Compute logits using the PLM
       $\mathcal{H}\leftarrow H(x\mid Y,f)$ // Calculate Entropy
       $x^{t}\leftarrow\text{Forward}(x_{0},\mathcal{H},t)$ // Forward at $t$
       $x^{t+1}\leftarrow\text{Forward}(x_{0},\mathcal{H},t+1)$ // Forward at $t+1$
       $L_{s}=-\sum_{i}q(x^{t}_{i}\mid x_{t+1})\log p_{\theta}(x^{t}_{i}\mid x_{t+1})$ // Loss calculation
@@
for $t=T$ to 00 do
       if $t=T$ then
             $x^{T}\leftarrow G_{M}(L,Y)$ // Initialize a sequence of $L$
       else
             $f_{\theta}\leftarrow M(x^{t},Y)$ // Compute logits at timestep $t$
@@
             $x^{t-1}\leftarrow\text{Sampling}(x^{t},l^{*},Y)$ // Sample from the previous timestep
@@
Algorithm 1 ENS and EAGS Algorithm
Our whole process is in Algorithm 1. Distinct from the prevailing methodologies in diffusion models as described by Ho et al. (2020a) and Austin et al. (2023), we do not employ the PLM parameterization strategy $\widetilde{p}_{\theta}(\widetilde{z}_{0}|z_{t},t)$, which preserves the original semantic embedding spaces during the training phase. We empirically find that PLM parameterization restricts the diversity of generated responses. Instead, to ensure the completeness of sentences, we implement Minimum Bayes Risk (MBR) decoding in the final generation step. We follow the traditional diffusion loss in Equation (Ho et al., 2020b), changing MSE with Cross Entrpy Loss.
@@ 4.1 Tasks and Datasets
Social Datasets The Paradetox dataset (Logacheva et al., 2022) focuses on removing profanities. The Deontology from ETHICS dataset (Hendrycks et al., 2023) evaluates ethical judgment based on scenarios. Paraphrase The QQP dataset (Wang et al., 2017) assesses paraphrase detection by treating one question as a paraphrase of another. Question Generation We use the Quasar-T dataset (Dhingra et al., 2017) to generate questions from passages and answers. Open-ended Generation The RocStories (Mostafazadeh et al., 2016) dataset is used to generate coherent story continuations from a given context. Detailed explanations are available in Appendix C and Dataset statistics are shown in Table 8.
@@ 4.2 Baselines
Since our objective is centered on generation tasks, we compare Diffusion-EAGS with four categories of baselines with the model similar to the size of RoBERTa-base (Liu et al., 2020): Auto-regressive Models (ARMs), CMLMs (Conditional Masked Language Models, Ghazvininejad et al. (2019a)), CDLMs, and DDLMs. For ARMs (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019), which is renowned for its proficiency in generation tasks. For CMLMs, we utilize CMLM (Ghazvininejad et al., 2019a) and DisCo (Kasai et al., 2020), which are transformer-based NAR models. Inference is carried out using the Mask-Predict (Ghazvininejad et al., 2019a) and the Easy-First (Kasai et al., 2020) inference algorithm respectively. For CDLMs, our baseline includes DiffuSeq (Gong et al., 2023a) and DINOISER (Ye et al., 2024). DiffuSeq is a diffusion model specifically designed for sequence-to-sequence applications, whereas DINOISER adaptively determines the range of sampled noise scales during training. For DDLMs, we utilize DiffusionBERT (He et al., 2022) , a state-of-the-art model in the research series of DiffusionLM (Li et al., 2022), and SEDD (Lou et al., 2024), a state-of-the-art model of open-ended generation in diffusion language generation domain. For SEDD, we download the pretrained version and fine-tune on it. Experimental results on LLMs can be found in Appendix E.